# Multi-State TD Target for Model-Free Reinforcement Learning

### v3

@@ VII Conclusion
We first rewrite (20):
$Q_{t+1}\left(\mathbf{s}_{t},\mathbf{a}_{t}\right)\leftarrow$ $(1-\alpha_{t})Q_{t}\left(\mathbf{s}_{t},\mathbf{a}_{t}\right)+\alpha_{t}\frac{1}{L}\sum_{l=1}^{L}$
$\sum_{i=1}^{l}\gamma^{i-1}\mathbf{r}_{t+i}+\max_{b\in\mathcal{A}}\gamma^{l}Q_{t}(\mathbf{s}_{t+l},\mathbf{b})$ (21)
Then we define
$\Delta_{t}(\mathbf{s}_{t})=Q_{t}(\mathbf{s}_{t},\mathbf{a}_{t})-Q^{*}(\mathbf{s}_{t},\mathbf{a}_{t})$ (22)
By subtracting $Q^{*}(\mathbf{s}_{t},\mathbf{a}_{t})$ from both side in (Proof.), we can obtain
$Q_{t+1}\left(\mathbf{s}_{t},\mathbf{a}_{t}\right)-Q^{*}(\mathbf{s}_{t},\mathbf{a}_{t})\leftarrow$ (23)
$\leftarrow(1-\alpha_{t})(Q_{t}\left(\mathbf{s}_{t},\mathbf{a}_{t}\right)-Q^{*}(\mathbf{s}_{t},\mathbf{a}_{t}))+\alpha_{t}\frac{1}{L}\sum_{l=1}^{L}\Big{(}\mathbf{r}_{l}$ (24)
$+\max_{b\in\mathcal{A}}\gamma^{l}Q_{t}(\mathbf{s}_{t+l},\mathbf{b})-Q^{*}(\mathbf{s}_{t},\mathbf{a}_{t})\Big{)},$ (25)
where $\mathbf{r}_{l}$ is $\sum_{i=1}^{l}\gamma^{i-1}\mathbf{r}_{t+i}$. According to (22), we have
$\Delta_{t+1}(\mathbf{s}_{t})$
$=$ $(1-\alpha_{t})\Delta_{t}(\mathbf{s}_{t})+\alpha_{t}\frac{1}{L}\sum_{l=1}^{L}\bigg{(}\mathbf{r}_{l}+\gamma^{l}\max_{b\in\mathcal{A}}Q_{t}(\mathbf{s}_{t+l},\mathbf{b})$
$-Q^{*}(\mathbf{s}_{t},\mathbf{a}_{t})\bigg{)}$
$=$ $(1-\alpha_{t})\Delta_{t}(\mathbf{s}_{t})+\alpha_{t}F_{t}(\mathbf{s}_{t},\mathbf{a}_{t}),$ (26)
where $F_{t}(\mathbf{s}_{t},\mathbf{a}_{t})$ is defined as:
$F_{t}(\mathbf{s}_{t},\mathbf{a}_{t})=$ $\frac{1}{L}\sum_{l=1}^{L}\bigg{(}\mathbf{r}_{l}+\gamma^{l}\max_{b\in\mathcal{A}}Q_{t}(\mathbf{s}_{t+l},\mathbf{b})-Q^{*}(\mathbf{s}_{t},\mathbf{a}_{t})\bigg{)}$ (27)
Now we want to proof $\alpha_{t}$ and $F_{t}(\mathbf{s}_{t},\mathbf{a}_{t})$ in (Proof.) meets the assumptions mentioned in Lemma 1. The first assumption is satisfied naturally as it is the same as the assumption in Theorem 1. For the assumption 2, we have
$\mathbb{E}\big{[}F_{t}(\mathbf{s}_{t},\mathbf{a}_{t})\big{]}=$ $\frac{1}{L}\sum_{l=1}^{L}\Bigg{(}\sum_{y\in\mathcal{X}}\mathrm{P}_{l}(\mathbf{s}_{t},y)\bigg{[}\mathbf{r}_{l}+\gamma^{l}\max_{b\in\mathcal{A}}Q_{t}(y,b)$
$-Q^{*}(\mathbf{s}_{t},\mathbf{a}_{t})\bigg{]}\Bigg{)}$ (28)
$=$ $\frac{1}{L}\sum_{l=1}^{L}(\mathbf{H}^{l}Q_{t})(\mathbf{s}_{t},\mathbf{a}_{t})-Q^{*}(\mathbf{s}_{t},\mathbf{a}_{t}),$ (29)
where the operator $\mathbf{H}^{l}$ is defined in Lemma 2. According to Lemma 2, we have
$\|\mathbb{E}\big{[}F_{t}(\mathbf{s}_{t},\mathbf{a}_{t})\big{]}\|_{\infty}=$ $\frac{1}{L}\sum_{i=1}^{L}(\mathbf{H}^{l}Q_{t})(\mathbf{s}_{t},\mathbf{a}_{t})-(\mathbf{H}^{l}Q^{*})(\mathbf{s}_{t},\mathbf{a}_{t})$
$\leq$ $\frac{1}{L}\sum_{l=1}^{L}\gamma^{l}\|Q_{t}-Q^{*}\|_{\infty}.$ (30)
Since $\gamma$ is smaller than 1, we have $\gamma^{l}\leq\gamma$, and results in
$\|\mathbb{E}\big{[}F_{t}(\mathbf{s}_{t},\mathbf{a}_{t})\big{]}\|_{\infty}$ $\leq\frac{1}{L}\sum_{l=1}^{L}\gamma^{l}\|Q_{t}-Q^{*}\|_{\infty}$
$\leq\frac{1}{L}\sum_{l=1}^{L}\gamma\|Q_{t}-Q^{*}\|_{\infty}$
$=\gamma\|Q_{t}-Q^{*}\|_{\infty}$
$=\|\Delta_{t}\|_{\infty}.$ (31)
The variance of $F_{t}(\mathbf{s}_{t},\mathbf{a}_{t})$ now can be calculated:
$\mathbf{var}$ $[F_{t}(\mathbf{s}_{t},\mathbf{a}_{t})]$
$=$ $\mathbb{E}\Bigg{[}\bigg{(}\frac{1}{L}\sum_{l=1}^{L}\big{(}\mathbf{r}_{l}+\gamma^{l}\max_{b\in\mathcal{A}}Q_{t}(\mathbf{s}_{t+l},\mathbf{b})-Q^{*}(\mathbf{s}_{t},\mathbf{a}_{t})\big{)}$
$-\frac{1}{L}\sum_{l=1}^{L}(\mathbf{H}^{l}Q_{t})(\mathbf{s}_{t},\mathbf{a}_{t})+Q^{*}(\mathbf{s}_{t},\mathbf{a}_{t})\bigg{)}^{2}\Bigg{]}$
$=$ $\mathbb{E}\Bigg{[}\bigg{(}\frac{1}{L}\sum_{l=1}^{L}\mathbf{r}_{l}+\gamma^{l}\max_{b\in\mathcal{A}}Q_{t}(\mathbf{s}_{t+l},\mathbf{b})$
$-\frac{1}{L}\sum_{l=1}^{L}(\mathbf{H}^{l}Q_{t})(\mathbf{s}_{t},\mathbf{a}_{t})\bigg{)}^{2}\Bigg{]}$
$=$ $\frac{1}{L^{2}}\mathbb{E}\Bigg{[}\bigg{(}\sum_{l=1}^{L}\mathbf{r}_{l}+\gamma^{l}\max_{b\in\mathcal{A}}Q_{t}(\mathbf{s}_{t+l},\mathbf{b})-\mathbb{E}\bigg{[}\sum_{l=1}^{L}\mathbf{r}_{l}$
$+\gamma^{l}\max_{b\in\mathcal{A}}Q_{t}(\mathbf{s}_{t+l},\mathbf{b})\bigg{]}\bigg{)}^{2}\Bigg{]}$
$=$ $\frac{1}{L^{2}}\mathbf{var}\bigg{(}\sum_{l=1}^{L}\gamma^{l}\max_{b\in\mathcal{A}}Q_{t}(\mathbf{s}_{t+l},\mathbf{b})\bigg{)}$
$=$ $\frac{1}{L^{2}}\mathbb{E}\bigg{[}\big{(}\sum_{l=1}^{L}\gamma^{l}\max_{b\in\mathcal{A}}Q_{t}(\mathbf{s}_{t+l},\mathbf{b})\big{)}^{2}\bigg{]}-\mathbb{E}\bigg{[}\sum_{l=1}^{L}\gamma^{l}$
$\max_{b\in\mathcal{A}}Q_{t}(\mathbf{s}_{t+l},\mathbf{b})\bigg{]}^{2}$
$\leq$ $\frac{1}{L^{2}}\mathbb{E}\bigg{[}\big{(}\sum_{l=1}^{L}\gamma^{l}\max_{b\in\mathcal{A}}Q_{t}(\mathbf{s}_{t+l},\mathbf{b})\big{)}^{2}\bigg{]}$
$\leq$ $\frac{1}{L^{2}}\bigg{(}\sum_{l=1}^{L}\gamma^{l}\max_{s\in\mathcal{S}}\max_{b\in\mathcal{A}}Q_{t}(\mathbf{s},\mathbf{b})\bigg{)}$
$\leq$ $\frac{1}{L^{2}}\bigg{(}\sum_{l=1}^{L}\gamma\max_{s\in\mathcal{S}}\max_{b\in\mathcal{A}}Q_{t}(\mathbf{s},\mathbf{b})\bigg{)}$
$\leq$ $\gamma^{2}\max_{s\in\mathcal{S}}\max_{b\in\mathcal{A}}(Q_{t}(\mathbf{s},\mathbf{b}))^{2}$
$\leq$ $\gamma^{2}\|\Delta_{t}+Q^{*}\|^{2}_{\infty}$
$=$ $\gamma^{2}\|\Delta_{t}\|_{\infty}^{2}+2\gamma^{2}\|\Delta_{t}\|_{\infty}\|Q^{*}\|_{\infty}+\gamma^{2}\|Q^{*}\|_{\infty}$ (32)
Notice that the reward function is bounded and $\|Q^{*}\|_{\infty}$ is non-negative and bounded, we discuss the convergence based on two situations:
  1. 1.
    If $\|\Delta_{t}\|_{\infty}\leq 1$:
    $\|\Delta_{t}\|_{\infty}^{2}+2\gamma^{2}\|\Delta_{t}\|_{\infty}\|Q^{*}\|_{\infty}+\gamma^{2}\|Q^{*}\|_{\infty}$
    $\leq\|\Delta_{t}\|_{\infty}^{2}+2\gamma^{2}\|Q^{*}\|_{\infty}+\gamma^{2}\|Q^{*}\|_{\infty}$ (33)
  2. 2.
    If $\|\Delta_{t}\|_{\infty}>1$, then $\|\Delta_{t}\|_{\infty}\leq\|\Delta_{t}\|_{\infty}^{2}$:
    $\|\Delta_{t}\|_{\infty}^{2}+2\gamma^{2}\|\Delta_{t}\|_{\infty}\|Q^{*}\|_{\infty}+\gamma^{2}\|Q^{*}\|_{\infty}$
    $\leq\gamma^{2}(1+2\|Q^{*}\|_{\infty})\|\Delta_{t}\|_{\infty}^{2}+\gamma^{2}\|Q^{*}\|_{\infty}^{2})$ (34)
We can choose
$C=\max(\gamma^{2}(1+2\|Q^{*}\|_{\infty}),2\gamma^{2}\|Q^{*}\|_{\infty}+\gamma^{2}\|Q^{*}\|^{2}_{\infty}),$ (35)
and we can obtain
$\mathbf{var}[F_{t}(\mathbf{s}_{t},\mathbf{a}_{t})]=C(1+\|\Delta_{t}\|_{\infty}^{2}),$ (36)
which satisfies the assumption 3 in Lemma 1.
According to Lemma 1, $\Delta_{t}$ converges to zero w.p.1, i.e., $Q_{t}(\mathbf{s},\mathbf{a})$ converges to $Q*(\mathbf{s},\mathbf{a})$ w.p.1. ∎
To this end, we prove that our multi-state TD algorithm can converge to the optimal Q-function.
Appendix B: Experiments Details
[…]
hyperparameters value
hidden width 256
$\tau$ 0.005
lr 0.0004
$\gamma$ 0.99
batch size 128
$\alpha$ 0.12
TABLE 2: Hyperparameters for SAC-related algorithms.
hyperparameters value
hidden width 256
$\tau$ 0.005
lr 0.0003
$\gamma$ 0.99
batch size 128
$\mu$ 0
$\theta$ 0.2
$\sigma$ 0.3
TABLE 3: Hyperparameters for DDPG-related algorithms.
Figure 3: Performance comparison of various DDPG algorithms in Walker-v4 environment.
Figure 4: Performance comparison of various SAC algorithms in Walker-v4 environment.
Figure 5: Performance comparison of various DDPG algorithms in HalfCheetah-v4 environment.
Figure 6: Performance comparison of various SAC algorithms in HalfCheetah-v4 environment.

### v4

@@ VII Conclusion
With $\operatorname{TD}^{\operatorname{error}}$ defined in (20), the equation (22) can be rewritten as:
$Q_{t+1}\left(\mathbf{s}_{t},\mathbf{a}_{t}\right)=$ $(1-\alpha_{t})Q_{t}\left(\mathbf{s}_{t},\mathbf{a}_{t}\right)+\alpha_{t}\frac{1}{L}\times$
$\sum_{l=1}^{L}\left(\bar{\mathbf{r}}_{t,l}+\max_{b\in\mathcal{A}}\gamma^{l}Q_{t}(\mathbf{s}_{t+l},b)\right).$ (23)
By subtracting $Q^{f}(\mathbf{s}_{t},\mathbf{a}_{t})$ from both side in (Proof.), we can obtain
$Q_{t+1}\left(\mathbf{s}_{t},\mathbf{a}_{t}\right)-Q^{f}(\mathbf{s}_{t},\mathbf{a}_{t})$
$=$ $(1-\alpha_{t})(Q_{t}\left(\mathbf{s}_{t},\mathbf{a}_{t}\right)-Q^{f}(\mathbf{s}_{t},\mathbf{a}_{t}))+\alpha_{t}\frac{1}{L}\times$
$\sum_{l=1}^{L}\left(\bar{\mathbf{r}}_{t,l}+\max_{b\in\mathcal{A}}\gamma^{l}Q_{t}(\mathbf{s}_{t+l},b)-Q^{f}(\mathbf{s}_{t},\mathbf{a}_{t})\right).$ (24)
Next, we define a function
$\Delta_{t}(s,a)=Q_{t}(s,a)-Q^{f}(s,a).$ (25)
As a result, (24) becomes
$\Delta_{t+1}(\mathbf{s}_{t},\mathbf{a}_{t})=$ $(1-\alpha_{t})\Delta_{t}(\mathbf{s}_{t},\mathbf{a}_{t})+\alpha_{t}F_{t}(\mathbf{s}_{t},\mathbf{a}_{t}),$ (26)
where $F_{t}(\mathbf{s}_{t},\mathbf{a}_{t})$ is defined as
$F_{t}(\mathbf{s}_{t},\mathbf{a}_{t})=$ $\frac{1}{L}\sum_{l=1}^{L}\left(\bar{\mathbf{r}}_{t,l}+\max_{b\in\mathcal{A}}\gamma^{l}Q_{t}(\mathbf{s}_{t+l},b)-Q^{f}(\mathbf{s}_{t},\mathbf{a}_{t})\right).$
We will apply Lemma 1 to (26) with $x=(\mathbf{s}_{t},\mathbf{a}_{t})$ to prove that the process converges to zero w.p.1, thereby completing the proof of the theorem. The first two assumptions of Lemma 1 are ensured by the corresponding assumptions in the theorem. What remains is to prove assumptions 3) and 4) of Lemma 1.
To prove assumption 3), we calculate
$\mathbb{E}\big{[}F_{t}(\mathbf{s}_{t},\mathbf{a}_{t})\big{]}=$ $\frac{1}{L}\sum_{l=1}^{L}\Bigg{(}\sum_{y\in\mathcal{S}}p_{l}(y|\mathbf{s}_{t},\mathbf{a}_{t})\bigg{[}\bar{\mathbf{r}}_{t,l}$
$+\gamma^{l}\max_{b\in\mathcal{A}}Q_{t}(y,b)-Q^{f}(\mathbf{s}_{t},\mathbf{a}_{t})\bigg{]}\Bigg{)}$
$=$ $\frac{1}{L}\sum_{l=1}^{L}\left[(\mathbf{H}^{l}Q_{t})(\mathbf{s}_{t},\mathbf{a}_{t})-(\mathbf{H}^{l}Q^{f})(\mathbf{s}_{t},\mathbf{a}_{t})\right],$
where the second equality follows from (19) and (21). Next, using the definition of the maximum norm and the property (18), we have
$\|\mathbb{E}\big{[}F_{t}(\mathbf{s}_{t},\mathbf{a}_{t})\big{]}\|_{\infty}\leq$ $\frac{1}{L}\sum_{i=1}^{L}\|\mathbf{H}^{l}Q_{t}-\mathbf{H}^{l}Q^{f}\|_{\infty}$
$\leq$ $\frac{1}{L}\sum_{l=1}^{L}\gamma^{l}\|Q_{t}-Q^{f}\|_{\infty}$ (27)
where $\gamma^{l}<\gamma<1$. This simplifies to
$\|\mathbb{E}\big{[}F_{t}(\mathbf{s}_{t},\mathbf{a}_{t})\big{]}\|_{\infty}$ $\leq\frac{1}{L}\sum_{l=1}^{L}\gamma\|Q_{t}-Q^{f}\|_{\infty}$
$=\gamma\|Q_{t}-Q^{f}\|_{\infty}=\|\Delta_{t}\|_{\infty}.$ (28)
This verifies assumption 3) of Lemma 1.
To prove assumption 4), we calculate the variance of $F_{t}(\mathbf{s}_{t},\mathbf{a}_{t})$ as follows:
$\operatorname{var}[F_{t}(\mathbf{s}_{t},\mathbf{a}_{t})]$
$=$ $\mathbb{E}\Bigg{[}\bigg{(}\frac{1}{L}\sum_{l=1}^{L}\big{(}\bar{\mathbf{r}}_{t,l}+\gamma^{l}\max_{b\in\mathcal{A}}Q_{t}(\mathbf{s}_{t+l},b)-Q^{f}(\mathbf{s}_{t},\mathbf{a}_{t})\big{)}$
$-\frac{1}{L}\sum_{l=1}^{L}(\mathbf{H}^{l}Q_{t})(\mathbf{s}_{t},\mathbf{a}_{t})+Q^{f}(\mathbf{s}_{t},\mathbf{a}_{t})\bigg{)}^{2}\Bigg{]}$
$=$ $\frac{1}{L^{2}}\mathbb{E}\Bigg{[}\bigg{(}\sum_{l=1}^{L}(\bar{\mathbf{r}}_{t,l}+\gamma^{l}\max_{b\in\mathcal{A}}Q_{t}(\mathbf{s}_{t+l},b))$
$-\sum_{l=1}^{L}(\mathbf{H}^{l}Q_{t})(\mathbf{s}_{t},\mathbf{a}_{t})\bigg{)}^{2}\Bigg{]}.$
Using the definition in (21), we have
$(\mathbf{H}^{l}Q_{t})(\mathbf{s}_{t},\mathbf{a}_{t})$ $=\mathbb{E}\bigg{[}\bar{\mathbf{r}}_{t,l}+\gamma^{l}\max_{b\in\mathcal{A}}Q_{t}(\mathbf{s}_{t+l},b)\bigg{]},$
which can simplify the variance expression as follows:
$\operatorname{var}[F_{t}(\mathbf{s}_{t},\mathbf{a}_{t})]$
$=$ $\frac{1}{L^{2}}\mathbb{E}\Bigg{[}\bigg{(}\sum_{l=1}^{L}(\bar{\mathbf{r}}_{t,l}+\gamma^{l}\max_{b\in\mathcal{A}}Q_{t}(\mathbf{s}_{t+l},b))$
$-\mathbb{E}\bigg{[}\sum_{l=1}^{L}(\bar{\mathbf{r}}_{t,l}+\gamma^{l}\max_{b\in\mathcal{A}}Q_{t}(\mathbf{s}_{t+l},b))\bigg{]}\bigg{)}^{2}\Bigg{]}$
$=$ $\frac{1}{L^{2}}\operatorname{var}\bigg{(}\sum_{l=1}^{L}\gamma^{l}(\bar{\mathbf{r}}_{t,l}+\max_{b\in\mathcal{A}}Q_{t}(\mathbf{s}_{t+l},b))\bigg{)}.$
Applying the variance property yields
$\operatorname{var}[F_{t}(\mathbf{s}_{t},\mathbf{a}_{t})]$
$\leq$ $\frac{1}{L^{2}}\mathbb{E}\bigg{[}\bigg{(}\sum_{l=1}^{L}\gamma^{l}(\bar{\mathbf{r}}_{t,l}+\max_{b\in\mathcal{A}}Q_{t}(\mathbf{s}_{t+l},b))\bigg{)}^{2}\bigg{]}$
$\leq$ $\frac{\gamma^{2}}{L^{2}}\mathbb{E}\bigg{[}\bigg{(}\sum_{l=1}^{L}(|\bar{\mathbf{r}}_{t,l}|+\max_{b\in\mathcal{A}}|Q^{f}(\mathbf{s}_{t+l},b)|$
$+\max_{b\in\mathcal{A}}|\Delta_{t}(\mathbf{s}_{t+l},b)|)\bigg{)}^{2}\bigg{]}$
$\leq$ $\gamma^{2}\bigg{(}\bar{r}_{\max}+\max_{s\in\mathcal{A},b\in\mathcal{A}}|Q^{f}(s,b)|+\max_{s\in\mathcal{A},b\in\mathcal{A}}|\Delta_{t}(s,b)|\bigg{)}^{2}$
$=$ $\gamma^{2}\bigg{(}C_{0}+\|\Delta_{t}\|_{\infty}\bigg{)}^{2}$
where $\bar{r}_{\max}$ is the maximum value of $\bar{\mathbf{r}}_{t,l}$, which is bounded because $\mathbf{r}_{t}$ is bounded, and $C_{0}=\bar{r}_{\max}+\max_{s\in\mathcal{A},b\in\mathcal{A}}|Q^{f}(s,b)|$.
If $C_{0}\leq 1$, then
$\operatorname{var}[F_{t}(\mathbf{s}_{t},\mathbf{a}_{t})]\leq\gamma^{2}\bigg{(}1+\|\Delta_{t}\|_{\infty}\bigg{)}^{2}.$
On the other hand, if $C_{0}>1$, we have
$\operatorname{var}[F_{t}(\mathbf{s}_{t},\mathbf{a}_{t})]\leq C_{0}^{2}\gamma^{2}\bigg{(}1+\|\Delta_{t}\|_{\infty}\bigg{)}^{2}.$
In both cases, assumption 4) of Lemma 1 is satisfied with $C=\gamma^{2}$ or $C=C_{0}^{2}\gamma^{2}$, respectively. Thus, the proof is complete. ∎